Applying Overhead to Jobs, Costing Jobs

Gorman Company builds internal conveyor equipment to client specifications. On October 1, Job 877 was in process with a cost of $18,690 to date.

During October, Jobs 878, 879, and 880 were started. Data on costs added during October for all jobs are as follows:

Job 877 Job 878 Job 879 Job 880
Direct materials $14,380 $5,970 $3,510 $1,770
Direct labor 14,600 8,500 2,200 2,400

Overhead is applied to production at the rate of 75% of direct labor cost. Job 877 was completed on October 28, and the client was billed at cost plus 50%. All other jobs remained in process.

Required:
1. Prepare a brief job-order cost sheet showing the October 1 balances of all four jobs, plus the direct materials and direct labor costs during October. There is no need to calculate applied overhead at this point or to total the costs.
2. Calculate the overhead applied to each job during October.
3. Calculate the balance in Work in Process on October 31.
4. What is the price of Job 877?
Answer #1

Part 1

Job 877

Job 878

Job 879

Job 880

Beginning balance

18690

Direct materials

14380

5970

3510

1770

Direct labor

14600

8500

2200

2400

Part 2

Applied overhead in October for:

Job 877 = $14,600 × 0.75 = $10,950

Job 878 = $8,500 × 0.75 = $6,375

Job 879 = $2,200 × 0.75 = $1,650

Job 880 = $2,400 × 0.75 = $1,800

Part 3

Work in Process, October 31:

Job 878 (5970+8500+6375)

$20845

Job 879 (3510+2200+1650)

$7360

Job 880 (1770+2400+1800)

$5970

Total

$34175

Part 4

Cost of Job 877 = 18690+14380+14600+10950 = $58620

Price of Job 877 = 58620+(58620*0.50) = $87930